# On the Scalability of GNNs for Molecular Graphs

### v3

@@ A.1 Pretraining
All models use 2-layer MLPs to encode node and edge features, respectively, followed by the core model of 16 layers of the MPNN++, Transformer or GPS++ (except for when scaling depth). To be able to tackle graph-level tasks, the outputs are aggregated to the graph level, e.g., by summing them up across the atoms of a molecule. Then, node and graph level representations go through separate 2-layer MLPs. Finally, representations are processed by separate task heads (2-layer MLPs) specific to each pretraining task. Further, all layers use layer norm and dropout with $p=0.1$. The encoder and model core additional have residual connections similar to the design in He et al. (2016a).
@@
Our hyperparameter search for all base models was conducted on all oberserved data samples with a constant model size of $10M\pm 0.1M$ parameters. For scaling on width, zero-shot scaling from $\mu P$ (Yang et al., 2022) was used. For other scaling results, $\mu P$ was used to scale the model with $10M$ parameters used as the base model. In the case of depth scaling, we adjusted the learning rate as suggested by depth-$\mu P$ (Yang et al., 2023). We did not consider adjusting the residual connections.
@@
Our base MPNN++, Transformer and hybrid GPS++ models are trained using Adam with a base learning rate of $0.003$, $0.001$, and $0.001$, respectively. We use 5 warm-up epochs followed by linear learning rate decay. All pretraining has been conducted with a batch size of 1024. Scaled version of the used models require advanced training strategies due to the large model size. We used multi-gpu training (with up to 8 NVIDIA A100-SXM4-40GB GPUs) and gradient accumulation, while adjusting batch size to keep the effective batch size constant.
A.2 Finetuning and Probing
@@
A.3 Performance Metrics
@@
Appendix B Trade-Off Between Over-smoothing and Depth
We note that GNN architectures exhibit over-smoothing phenomenon, which implies that latent representations of a network become similar and coarser as the network grows in size. Prior evidence suggests that over-smoothing occurs linearly with the increasing depth of GNN networks (Hamilton, 2020; Xu et al., 2019). We observed similar behaviors for MPNN architectures during pretraining where the performance for node-level tasks degrades significantly with very deep networks. However, it is difficult to determine without any doubt that over-smoothing is the culprit.
On another hand, over-smoothing is believed to be aleviated by graph Transformers. Recent works argue that Transformers present favorable properties which make them robust towards over-smoothing, such as the provision of embeddings and the inductive bias of attention (Dovonon et al., 2024). However, we still observe a degradation of performance with depth of our Transformer models, in contradiction with this hypothesis. Its theoretical understanding and empirical analysis remains an open question for future work.
Appendix C Additional Related Work
C.1 Foundational Models for Molecules
Here, we present additional advancements in foundational models making use of molecular graphs. Recent works have argued that the use of high-capacity models will be a significant boon to scientific discovery tasks (Wang et al., 2023a). Of specific interest are tasks in the quantum and molecular discovery paradigms (Zhang et al., 2023) which demand domain-specific expertise such as knowledge of structure, provision of additional inductive biases and large data requirements. Towards this hypothesis, Fifty et al. (2023) present an in-context learning framework for molecular property prediction without explicitly using a meta learning procedure. This leads to a general algorithm capable of discovering high-level structures from a pretraining sample set. Guo et al. (2021) propose a similar framework making use of few-shot learning techniques resulting in a sample-efficient learning procedure. Xu et al. (2023) present an alternative approach by modelling the full graphical structure of molecules across different property prediction tasks. Although effective, modelling the entire graph results in a computationally intensive learning procedure. Finally, Cao et al. (2023) scale up learning to larger graph sizes by segmenting graph neighborhoods on the fly. An ad-hoc partitioning procedure is employed and interleaved with the learning phase in order to accelerate learning on larger and dense graphical clusters.
C.2 Expressivity of GNNs
Prior work highlights that GNN architectures are limited in their expressivity to distinguish between graphs of similar node arrangements but different geometrical structures (Xu et al., 2019). Various works indicate this as a consequence of aggregation functions and other design factors involved in GNN training (Hamilton, 2020). On the other hand, recent work argues that only specific architectures are found robust to over-smoothing when building latent representations (Dovonon et al., 2024). For instance, graph Transformers exhibit over-smoothing robustness as they utilize strong inductive biases such as attention. Xu et al. (2021a) connect the limited expressivity of GNNs with their ability to extrapolate on simpler tasks. Contrary to multi-layer networks, GNNs struggle to extrapolate on simpler tasks but show promise for improvement. Morris et al. (2019) aim to tackle over-smoothing by building higher-order GNN architectures capable of capturing intricate node characteristics in their deeper layers. Finally, Ying et al. (2018) present the differentiable pooling module capable of pooling neighboring node features which aid in reducing noise across layer representations.
@@ E.1 Width Scaling
Figure 9: Comparison of probing and finetuning for MPNN++ (left), Transformer (center), and hybrid GPS++ (right) across different model sizes on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlations for MPNN++, Transformer and GPS++ models show improving scaling behavior with increasing number of parameters across the TDC benchmark. The average Spearman correlation between width and performance for probing is 0.69, 0.82 and 0.73, respectively, and 0.72 when finetuning MPNN++, effectively showing that model size plays an important role in predictive performance.
@@ E.2 Depth Scaling
Figure 10: Comparison of probing for different model depths on the Polaris benchmark with MPNN++ (left), Transformer (center), and hybrid GPS++ (right). Darker green shades denote higher/desirable metric values. Average Spearman correlation between depth and performance is $0.47$, $0.55$, and $0.50$, respectively. Probing shows positive scaling trend with increasing depth across the Polaris benchmark.
@@
Figure 11: Comparison of probing and finetuning for MPNN++ (left), Transformer (center), and GPS++ (right) models across different model depths on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between depth and probing performance is -0.11, 0.27 and -0.03, respectively, and 0.33 for finetuned MPNN++. While performance mostly increases with depths up to 8, larger depths often show signs of saturation.
@@ E.3 Molecule Scaling
Refer to captionRefer to captionRefer to caption
Figure 12: Scaling behavior of probed 100M parameter models across different dataset molecule fractions on the Polaris benchmark with MPNN++ (left), Transformer (center), and hybrid GPS++ (right). Darker green shades denote higher/desirable metric values. Average Spearman correlation between molecule fraction and probing performance is 0.30, 0.73, and 0.40, respectively. Models show consistent improvement in performance with the increasing size of datasets.
@@
Figure 13: Comparison of probing and finetuning for MPNN++ (left), Transformer (center) and GPS++ (right) models across different dataset sizes on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between molecule fraction and probing performance is 0.28, 0.13 and 0.15, respectively, and 0.32 for finetuned MPNN++. Finetuned models scale better when compared to probed models. However, increasing the size of finetuning datasets leads to minor improvements beyond the 50% dataset size fraction.
@@ E.4 Label Scaling
Figure 14: Performance of MPNN++ probed models across different label fractions in the Polaris benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between label fraction and probing performance is 0.57.
Figure 15: Comparison of MPNN++ probing and finetuning across different label fractions on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between label fraction and performance is 0.54 for probed MPNN++ and 0.37 for finetuned MPNN++. Finetuned models scale better when compared to probed models. Increasing label fractions do not deteriorate model performance.
@@ E.5 Dataset Ablation
Figure 16: Comparison of probed 100M parameter MPNN++ models on the Polaris benchmark tasks (in columns) after pretrained without certain pretraining datasets (in rows). Darker green shades denote higher/desirable metric values. We observe that removing PCBA_1328 significantly hurts downstream performance across almost all tasks, while removing the L1000 leads to noticeable improvements on most tasks.
Figure 17: Performance of probed 100M parameter MPNN++ models on TDC benchmark tasks (in columns) after pretrained without certain datasets (in rows). Darker green shades denote higher/desirable metric values. We observe that removing PCBA_1328 significantly hurts downstream performance across almost all tasks, while removing the L1000 leads to noticeable improvements on most tasks.
@@ E.6 Task-Head Ablation
Figure 18: Performance of probed 100M parameter MPNN++ models on TDC benchmark (left) and Polaris benchmark (right), with the tasks in columns, when probing from different task heads (in rows). Darker green shades denote higher/desirable metric values, and bold/underline indicates the best value in a given column. We observe that the graph_output_nn (e.g. the hidden representation that is fed to all task-heads) is overall the best choice for probing. We hypothesize this is because it captures and compresses the combined information from the various pretraining tasks. The PCBA_1328 task-head is also a good choice due to its proximity to the considered downstream tasks. The PCQM4M_G25 task-head is the least useful as the tasks are fundamentally different from the downstream tasks.
@@ E.7 The TDC Benchmark – Data Leakage
PCBA_1328 contains only classification assays with more than 6000 molecules each, which automatically disqualifies most of TDC and all of Polaris. The TDC datasets remaining after this constraint are the 3 CYP*_Veith datasets, and the AMES dataset. The AMES dataset is not present in PubChem (Kim et al., 2023), excluding it from the list of potential leaks. Regarding the 3 CYP*_Veith datasets, they represent inibition assays against recombinant enzymes (Veith et al., 2009). The three assays from TDC, and two others from the paper, are all present and aggregated under assayID-1851. Therefore, whenever a molecule is active against any of the enzyme, the value is $1$, otherwise it is 00. Therefore, there is a minor leak, although the datasets are not identical. Further, no evidence of leak was observed in terms of abnormally high performance of the model on these assays, which is expected considering that the model is learning more than 3000 labels simultaneously.
Appendix F Scaling Law Details
@@
Table 1 presents values of $\alpha$ (Equation 1) based on model parameters and tasks considered in our downstream task experiments on the Polaris benchmark (Figure 3). We choose $|\theta_{c}|=1B$ and fix our final training error values corresponding to this model’s performance. On average, we see $\alpha\approx 0.081$ for probing and $\alpha\approx 0.098$ for finetuning. These relationships hold across 6 orders of magnitude in $|\theta|$ indicating that finetuning behavior scales logarithmically with the number of trainable parameters. It is also worth noting that Kaplan et al. (2020) obtain similar $\alpha\approx 0.076$ indicating that our power law fit lies within the same parameter budget.
Table 2 compares values of $\beta$ (Equation 2) for pretraining our models as presented in Figure 2. We choose $|\mathcal{D}_{c}|=5M$ to be $100\%$ of molecular data. As observed, $\beta\approx 0.110$ for MPNN++, $\beta\approx 0.106$ for Transformer and $\beta\approx 0.106$ for GPS++ for the overall test error. This holds for datasets up to 5M molecules. Notably, all $\beta>0$ indicate that, given a computational budget, larger GNNs scale favorably. We again note that Kaplan et al. (2020) have similar power law fits of $\beta\approx 0.095$ albeit with a $10^{13}$ token corpus. This allows us to draw two conclusions. Firstly, GNNs continue to demonstrate optimal scaling with limited datasets. Secondly, our 5M molecules (10M node features) dataset is sufficient to demonstrate scaling behavior equivalent to a $10^{13}$ token language corpus, up to the 1B parameters regime.

### v4

@@ B.1 Pretraining
All models use 2-layer MLPs to encode node and edge features, respectively, followed by the core model of 16 layers of the MPNN++, Transformer or GPS++ (except for when scaling depth). To be able to tackle graph-level tasks, the outputs are aggregated to the graph level, e.g., by summing them up across the atoms of a molecule. Then, node and graph level representations go through separate 2-layer MLPs. Finally, representations are processed by separate task heads (2-layer MLPs) specific to each pretraining task. Further, all layers use layer norm and dropout with $p=0.1$. The encoder and model core additional have residual connections similar to the design in He et al. [22].
@@
Our hyperparameter search for all base models was conducted on all oberserved data samples with a constant model size of $10M\pm 0.1M$ parameters. For scaling on width, zero-shot scaling from $\mu P$ [76] was used. For other scaling results, $\mu P$ was used to scale the model with $10M$ parameters used as the base model. In the case of depth scaling, we adjusted the learning rate as suggested by depth-$\mu P$ [77]. We did not consider adjusting the residual connections.
@@
Our base MPNN++, Transformer and hybrid GPS++ models are trained using Adam with a base learning rate of $0.003$, $0.001$, and $0.001$, respectively. We use 5 warm-up epochs followed by linear learning rate decay. All pretraining has been conducted with a batch size of 1024. Scaled version of the used models require advanced training strategies due to the large model size. We used multi-gpu training (with up to 8 NVIDIA A100-SXM4-40GB GPUs) and gradient accumulation, while adjusting batch size to keep the effective batch size constant. Most models were trained on sigle gpus but our 300M and 1B parameter models used 4 and 8 gpus, respectively.
B.2 Finetuning and Probing
@@
We finally note that all downstream task experiments were conducted on single cpus.
B.3 Performance Metrics
@@
[…]
Appendix C Trade-Off Between Over-smoothing and Depth
We note that GNN architectures exhibit over-smoothing phenomenon, which implies that latent representations of a network become similar and coarser as the network grows in size. Prior evidence suggests that over-smoothing occurs linearly with the increasing depth of GNN networks [21, 72]. We observed similar behaviors for MPNN architectures during pretraining where the performance for node-level tasks degrades significantly with very deep networks. However, it is difficult to determine without any doubt that over-smoothing is the culprit.
On another hand, over-smoothing is believed to be aleviated by graph Transformers. Recent works argue that Transformers present favorable properties which make them robust towards over-smoothing, such as the provision of embeddings and the inductive bias of attention [14]. However, we still observe a degradation of performance with depth of our Transformer models, in contradiction with this hypothesis. Its theoretical understanding and empirical analysis remains an open question for future work.
@@ E.1 Width Scaling
Figure 9: Width Scaling: Comparison of probing and finetuning for MPNN++ (left), Transformer (center), and hybrid GPS++ (right) across different model sizes on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlations for MPNN++, Transformer and GPS++ models show improving scaling behavior with increasing number of parameters across the TDC benchmark. The average Spearman correlation between width and performance for probing is 0.69, 0.82 and 0.73, respectively, and 0.72 when finetuning MPNN++, effectively showing that model size plays an important role in predictive performance.
@@ E.2 Depth Scaling
Figure 10: Depth Scaling: Comparison of probing for different model depths on the Polaris benchmark with MPNN++ (left), Transformer (center), and hybrid GPS++ (right). Darker green shades denote higher/desirable metric values. Average Spearman correlation between depth and performance is $0.47$, $0.55$, and $0.50$, respectively. Probing shows positive scaling trend with increasing depth across the Polaris benchmark.
@@
Figure 11: Depth Scaling: Comparison of probing and finetuning for MPNN++ (left), Transformer (center), and GPS++ (right) models across different model depths on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between depth and probing performance is -0.11, 0.27 and -0.03, respectively, and 0.33 for finetuned MPNN++. While performance mostly increases with depths up to 8, larger depths often show signs of saturation.
@@ E.3 Molecule Scaling
Figure 12: Molecule Scaling: Scaling behavior of probed 100M parameter models across different dataset molecule fractions on the Polaris benchmark with MPNN++ (left), Transformer (center), and hybrid GPS++ (right). Darker green shades denote higher/desirable metric values. Average Spearman correlation between molecule fraction and probing performance is 0.30, 0.73, and 0.40, respectively. Models show consistent improvement in performance with the increasing size of datasets.
Refer to caption Refer to caption
Figure 13: Molecule Scaling: Comparison of probing and finetuning for MPNN++ (left), Transformer (center) and GPS++ (right) models across different dataset sizes on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between molecule fraction and probing performance is 0.28, 0.13 and 0.15, respectively, and 0.32 for finetuned MPNN++. Finetuned models scale better when compared to probed models. However, increasing the size of finetuning datasets leads to minor improvements beyond the 50% dataset size fraction.
@@ E.4 Label Scaling
Figure 14: Label Scaling: Performance of MPNN++ probed models across different label fractions in the Polaris benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between label fraction and probing performance is 0.57.
Figure 15: Label Scaling: Comparison of MPNN++ probing and finetuning across different label fractions on the TDC benchmark. Darker green shades denote higher/desirable metric values. Average Spearman correlation between label fraction and performance is 0.54 for probed MPNN++ and 0.37 for finetuned MPNN++. Finetuned models scale better when compared to probed models. Increasing label fractions do not deteriorate model performance.
@@ E.5 Dataset Ablation
Figure 16: Dataset Ablation: Comparison of probed 100M parameter MPNN++ models on the Polaris benchmark tasks (in columns) after pretrained without certain pretraining datasets (in rows). Darker green shades denote higher/desirable metric values. We observe that removing PCBA_1328 significantly hurts downstream performance across almost all tasks, while removing the L1000 leads to noticeable improvements on most tasks.
Figure 17: Dataset Ablation: Performance of probed 100M parameter MPNN++ models on TDC benchmark tasks (in columns) after pretrained without certain datasets (in rows). Darker green shades denote higher/desirable metric values. We observe that removing PCBA_1328 significantly hurts downstream performance across almost all tasks, while removing the L1000 leads to noticeable improvements on most tasks.
@@ E.6 Task-Head Ablation
Figure 18: Task-Head Ablation: Performance of probed 100M parameter MPNN++ models on TDC benchmark (left) and Polaris benchmark (right), with the tasks in columns, when probing from different task heads (in rows). Darker green shades denote higher/desirable metric values, and bold/underline indicates the best value in a given column. We observe that the graph_output_nn (e.g. the hidden representation that is fed to all task-heads) is overall the best choice for probing. We hypothesize this is because it captures and compresses the combined information from the various pretraining tasks. The PCBA_1328 task-head is also a good choice due to its proximity to the considered downstream tasks. The PCQM4M_G25 task-head is the least useful as the tasks are fundamentally different from the downstream tasks.
@@ E.7 The TDC Benchmark – Data Leakage
PCBA_1328 contains only classification assays with more than 6000 molecules each, which automatically disqualifies most of TDC and all of Polaris. The TDC datasets remaining after this constraint are the 3 CYP*_Veith datasets, and the AMES dataset. The AMES dataset is not present in PubChem [29], excluding it from the list of potential leaks. Regarding the 3 CYP*_Veith datasets, they represent inibition assays against recombinant enzymes [66]. The three assays from TDC, and two others from the paper, are all present and aggregated under assayID-1851. Therefore, whenever a molecule is active against any of the enzyme, the value is $1$, otherwise it is 00. Therefore, there is a minor leak, although the datasets are not identical. Further, no evidence of leak was observed in terms of abnormally high performance of the model on these assays, which is expected considering that the model is learning more than 3000 labels simultaneously.
Appendix F Comparison of MolGPS to Unsupervised Methods
Table 1: Comparison of MolGPS variants to the self-supervised GraphMVP model on the MoleculeNet dataset (test AUROC). Note that we only consider the datasets that were not part of our pretraining.
Method BACE BBBP Clintox Sider
GraphMVP 0.812 0.724 0.775 0.639
1B MolGPS w/o Phenomics 0.806 0.802 0.797 0.649
1B MolGP 0.828 0.8 0.809 0.67
3B MolGPS 0.832 0.809 0.807 0.666
Compared to unsupervised pretraining approaches, it is important to remark that the scale of pretraining data is not directly comparable to that of our supervised pretraining approach, i.e, billions of molecules in some cases. We note that previous works in the GNN literature [12] found interesting scaling trends despite scaling to even less than the 5M graphs that we have aggregated in the LargeMix dataset.
In our supervised context, the data scale does not only depend on the number of molecules seen during pretraining. Instead, each molecule should be considered in conjunction with the set of pretraining labels. PCBA_1328 considers more than 1k different labels per molecule (albeit with high sparsity) and PCQM4M comes with 25 graph-level tasks and 4 node-level tasks (for each node of the  4M molecules). We point to our label scaling study that shows the impact of reduced data diversity by removing labels (e.g., Figure 2) for convincing evidence for this trend. The performance gains from incorporating Phenomics data into the pretraining data mix also support this claim, where the addition of  500k molecules with  6k highly informative labels per graph (Figure 5) leads to strong performance gains.
Here, we provide further comparison to unsupervised pretraining approaches (Table 1 and 2). We observe that the unsupervised MolE model variants [43] clearly underperform the standard MolE model that leverages both unsupervised and supervised pretraining (Table 2), which is in turn outperformed by the listed MolGPS variants by large margins. Even our smaller 100M MPNN++ models (that are of comparable parametric size to MolE) both outperform the self-supervised variants (see Figure 5).
In Table 1, we compare to the self-supervised GraphMVP [35] model on MoleculeNet. MolGPS without Phenomics pretraining outperforms GraphMVP on 3/4 tasks, while the model variants that were pretrained with additional Phenomics data outperform across all tasks.
Table 2: Comparison of MolGPS variants to the self-supervised variants of MolE and the standard MolE (supervised + self-supervised pretraining) on TDC benchmark collection. Supervised pretraining significantly improves MolE performance as seen in the normalized performance (left-most column). MolGPS models outperform all MolE models by a large margin, exhibiting the best performance in all but two tasks.
Lipophilicity_AstraZeneca
Caco2_Wang
LD50_Zhu
Solubility_AqSolDB
PPBR_AZ
BBB_Martins
HIA_Hou
Pgp_Broccatelli
Bioavailability_Ma
CYP3A4_Substrate_CarbonMangels
AMES
hERG
DILI
VDss_Lombardo
Half_Life_Obach
Clearance_Microsome_AZ
Clearance_Hepatocyte_AZ
CYP2D6_Substrate_CarbonMangels
CYP2C9_Substrate_CarbonMangels
CYP2D6_Veith
CYP3A4_Veith
CYP2C9_Veith
Method
Norm.
MAE
MAE
MAE
MAE
MAE
AUROC
AUROC
AUROC
AUROC
AUROC
AUROC
AUROC
AUROC
Spear.
Spear.
Spear.
Spear.
AUPRC
AUPRC
AUPRC
AUPRC
AUPRC
MolE-FuncEnv (unsupervised) 0.47 0.46 0.355 0.597 0.799 8.57 0.895 0.951 0.873 0.638 0.612 0.831 0.871 0.89 0.622 0.579 0.567 0.373 0.715 0.411 0.678 0.857 0.759
MolE-AtomEnv (unsupvervised) 0.477 0.464 0.471 0.582 0.81 8.191 0.895 0.949 0.871 0.683 0.633 0.832 0.844 0.883 0.596 0.518 0.531 0.367 0.706 0.429 0.665 0.865 0.773
MolE 0.728 0.469 0.31 0.577 0.792 8.073 0.903 0.963 0.915 0.654 0.67 0.813 0.823 0.883 0.654 0.549 0.607 0.381 0.699 0.446 0.682 0.867 0.801
1B MolGPS w/o Phenomics 1.222 0.4 0.347 0.645 0.714 6.249 0.922 0.984 0.941 0.64 0.681 0.8389 0.86 0.937 0.655 0.64 0.659 0.56 0.712 0.483 0.747 0.905 0.871
1B MolGPS 1.305 0.391 0.288 0.589 0.706 6.497 0.939 0.975 0.947 0.686 0.681 0.85 0.868 0.933 0.649 0.632 0.649 0.527 0.700 0.474 0.741 0.898 0.832
3B MolGPS 1.404 0.386 0.292 0.557 0.679 6.464 0.941 0.98 0.948 0.701 0.68 0.857 0.864 0.942 0.649 0.631 0.633 0.57 0.713 0.464 0.75 0.9 0.838
Appendix G Scaling Law Details
@@
Table 3 presents values of $\alpha$ (Equation 1) based on model parameters and tasks considered in our downstream task experiments on the Polaris benchmark (Figure 3). We choose $|\theta_{c}|=1B$ and fix our final training error values corresponding to this model’s performance. On average, we see $\alpha\approx 0.081$ for probing and $\alpha\approx 0.098$ for finetuning. These relationships hold across 6 orders of magnitude in $|\theta|$ indicating that finetuning behavior scales logarithmically with the number of trainable parameters. It is also worth noting that Kaplan et al. [28] obtain similar $\alpha\approx 0.076$ indicating that our power law fit lies within the same parameter budget.
Table 4 compares values of $\beta$ (Equation 2) for pretraining our models as presented in Figure 2. We choose $|\mathcal{D}_{c}|=5M$ to be $100\%$ of molecular data. As observed, $\beta\approx 0.110$ for MPNN++, $\beta\approx 0.106$ for Transformer and $\beta\approx 0.106$ for GPS++ for the overall test error. This holds for datasets up to 5M molecules. Notably, all $\beta>0$ indicate that, given a computational budget, larger GNNs scale favorably. We again note that Kaplan et al. [28] have similar power law fits of $\beta\approx 0.095$ albeit with a $10^{13}$ token corpus. This allows us to draw two conclusions. Firstly, GNNs continue to demonstrate optimal scaling with limited datasets. Secondly, our 5M molecules (10M node features) dataset is sufficient to demonstrate scaling behavior equivalent to a $10^{13}$ token language corpus, up to the 1B parameters regime.